OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x - 7 / x + 2 and g(x) = -2x-7 / x-1

OpenStudy (anonymous):

Please help :) I forgot how to do this but you should have covered it already.

OpenStudy (anonymous):

yes i did :) hold on

OpenStudy (anonymous):

Oh yay you're typing i'm still here

OpenStudy (anonymous):



OpenStudy (anonymous):

its gonn start like tht

OpenStudy (anonymous):

I did get that far. My big question is after this step and how to take it. Do I flip the denominator and multiply both terms?

OpenStudy (anonymous):

\[f(x) =\frac{ x - 7 }{ x + 2}\] \[g(x)=\frac{-2x-7}{x-1}\]

OpenStudy (anonymous):

what @kiimiilee wrote is correct here is the first step :

OpenStudy (anonymous):

\[f(g(x))=f(\frac{-2x-7}{x-1})=\frac{ \frac{-2x-7}{x-1}- 7 }{ \frac{-2x-7}{x-1} + 2}\]

OpenStudy (anonymous):

Does x=-7/2?

OpenStudy (anonymous):

from there on in it is a bunch of annoying algebra, which will lead to a whole raft of cancellation, and you should end up only with \(x\) at the end

OpenStudy (anonymous):

no, \(x\) is not a number at all your job is to compose the function and find that you end up with the identity, i.e. that \[f(g(x))=x\]

OpenStudy (anonymous):

So if I demonstrate that both f(g(x)) and g(f(x)) = -7/2 would I be answering the questioN?

OpenStudy (anonymous):

yes but it would be wrong

OpenStudy (anonymous):

Ok then where would I go from there.

OpenStudy (anonymous):

you simplify i think @satellite73

OpenStudy (anonymous):

multiply top and bottom by x-1

OpenStudy (anonymous):

OK then what?

OpenStudy (anonymous):

you keep canceling and you'll get x

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

np :)